A diffraction grating with N = 8000 lines per cm deviates light with λ = 300 nm. What is the highest order that can be observed? The answer must be an integer.
Answer #1

The diffraction condition (e+d)sin\theta =n\lambda where e+d= grating element

he conditionfor the highest order   

nmax = \frac{(e+d) }{\lambda }

= 1 cm / 8000 * 300 * 10^-7 cm.

= 4.16 so the answer is 4